# Tony Lindeberg

### v1

@@ 3.3 Visualizations of spatial and spatio-temporal receptive fields
Figures 13 show sample illustrations of receptive fields according to the generalized Gaussian derivative model for visual receptive fields. Figures 1 shows purely spatial receptive fields of the form
$T_{\varphi^{m}}(x;\;s,\Sigma)=\partial_{\varphi^{m}}g(x;\;s,\Sigma)$ (14)
@@
$\begin{split}\lambda_{2}&=\sigma_{2}^{2},\end{split}$ (19)
which then leads to the following explicit expression for the affine Gaussian derivative kernel (Lindeberg 2024 Equations (163) and (164))
@@ 3.4 Relations to scale-normalized derivatives
In Lindeberg (2021; 2025b; 2025c), so-called unnormalized spatio-temporal derivative expressions of the above forms are additionally combined with different forms of scale normalization, as depending on the actual values of the spatial scale parameter $s$, the spatial covariance matrix $\Sigma$ and the temporal scale parameter $\tau$, to define scale-normalized spatial and spatio-temporal derivative operators [^2] For in-depth treatments about how to extend the regular unnormalized spatial and spatio-temporal derivative operators ${\cal D}_{x}$ and ${\cal D}_{x,t}$ to corresponding scale-normalized spatial and spatio-temporal derivative operators, that make the full derivative-based receptive field responses in (7) and (8) provably covariant under the composed geometric image transformations in (25), (32), (33) and (34), see Lindeberg (2025b; 2025c). Basically, to achieve covariance under with respect to uniform spatial scaling transformations of the form $x^{\prime}=S_{x}\,x$, where $S_{x}\in{\mathbb{R}}_{+}$ is a spatial scaling factor, the partial derivative operators $\partial_{x_{i}}$ in the spatial gradient operator $\nabla_{x}$ according to (10) and the spatial Hessian operator ${\cal H}_{x}$ according to (11) should be replaced by regular scale-normalized spatial derivatives according to $\partial_{x_{i},\text{norm}}=s^{1/2}\,\partial_{x_{i}}$ according to Lindeberg (1998) Equation (6) for the scale normalization power $\gamma=1$, while the regular directional derivative operator $\partial_{\varphi}^{m}$ according to (9) should be replaced by the scale-normalized directional derivative operator $\partial_{\varphi,\text{norm}}^{m}$ according to Lindeberg (2025b) Equation (33). To achieve covariance with respect to non-isotropic spatial affine transformations of the form $x^{\prime}=A\,x$, where $A$ is a positive definite $2\times 2$ affine transformation matrix, the regular gradient operator $\nabla_{x}$ according to (10) should be replaced by the scale-normalized affine gradient operator $\nabla_{x,\text{affnorm}}$ according to Lindeberg (2025b) Equation (111), while the regular Hessian operator ${\cal H}_{x}$ according to (11) should be replaced by the scale-normalized affine Hessian operator ${\cal H}_{x,\text{affnorm}}$ according to Lindeberg (2025b) Equation (140). To achieve both covariance under Galilean transformations of the form $x^{\prime}=x+u\,t$, where $u=(u_{1},u_{2})^{T}\in{\mathbb{R}}^{2}$ is a 2-D motion vector, and covariance under temporal scaling transformations of the form $t^{\prime}=S_{t}\,t$, where $S_{t}\in{\mathbb{R}}_{+}$ is a temporal scaling factor, the temporal differentiation of the receptive fields should be performed in terms of scale-normalized velocity-adapted temporal derivative operators obtained by replacing the regular velocity-adapted temporal derivative operator $\partial_{\bar{t}}^{n}$ in (12) by its scale-normalized correspondence $\partial_{\bar{t},\text{norm}}^{n}$ according to Lindeberg (2025b) Equation (168). [^3] If using scale-normalized derivative operators in the models for the visual receptive fields, it should, however, be observed that the introduction of scale-normalized derivatives will substantially influence the evolution properties over scales of the receptive field responses. This follows, since the scale-dependent scale normalization factors used in the formulations of scale-normalized derivative operators, as extensively described in Lindeberg (2025b), would then also need to be differentiated with respect to the spatial and the temporal scale parameters. For this reason, we do in this paper isolate the evolution properties over the filter parameters to only receptive field models expressed in terms of regular, not scale-normalized, spatial and/or spatio-temporal derivative expressions. If the use of scale-normalized derivatives is warranted, as it can be because of the requirement of making it possible to match the output of derivative-based receptive field operators on the image data, this can be easily accomplished by performing the scale normalization in a second stage, after the computation of regular, not scale-normalized, derivative operators in a first processing stage. that are provably scale covariant, in the sense that the magnitude of the filter responses can be perfectly matched under specific families of geometric image transformations. In this treatment, we will, however, not consider the effects of such scale normalization, and will instead focus solely on the effects of the either purely spatial or joint spatio-temporal smoothing operations obtained by convolution with kernels of the forms (1) or (2) in combination with either purely spatial or joint spatio-temporal derivative computations, in turn leading to receptive field responses of the forms (7) or (8). For the task of relating the responses of linear receptive fields of the forms (7) or (8), such an analysis is sufficient, provided that the spatial differentiation operators ${\cal D}_{x}$ and the spatio-temporal differential operators ${\cal D}_{x,t}$ in these receptive field models are purely linear operators.
@@ 3.5 Covariance properties of the spatial and the spatio-temporal smoothing kernels
With regard to variabilities of the shapes of spatial or spatio-temporal receptive fields in relation to variabilities in spatial or spatio-temporal image transformations involved in the image formation process that determines the structures of the either purely spatial or joint spatio-temporal image data $f$, an important consequence of the formulation of the spatial smoothing kernel $T(x;\;s,\Sigma)$ according to (1) and the formulation of the joint spatio-temporal smoothing kernel $T(x,t;\;s,\Sigma,\tau,v)$ according to (2) is that these forms of convolution kernels are covariant under important basic classes of geometric image transformations.
Figure 4: Illustration of the geometry underlying the composed locally linearized projection models in Equation (25) and Equations (32)–(34), when extended to a multi-view imaging situation, with each view indexed by $k$. We consider a local, in the spatio-temporal case possibly moving, surface patch, which is projected to an arbitrary view in a multi-view locally linearized projection model, with the fixation point $F$ on the surface mapped to the origin $O^{(k)}=0$ in the image plane for the observer with the optic center $P^{(k)}$. Then, any point in the tangent plane to the surface at the fixation point, as parameterized by the local coordinates $\xi$ in a coordinate frame attached to the tangent plane of the surface with $\xi=0$ at the fixation point $F$, is by the local linearization mapped to the image point $x^{(k)}$. (Figure reproduced from Lindeberg (2025b) with permission (OpenAccess).)
3.5.1 Purely spatial covariance property
@@
$f^{\prime}(x^{\prime})=f(x)\quad\quad\mbox{for}\quad\quad x^{\prime}=S_{x}\,A$ (25)
@@
provided that the parameters of the receptive fields are matched according to Lindeberg (2025b) Equations (252)–(253):
@@
Due to the overparameterization of the degrees of freedom spanned by the spatial scale parameter $s$ and the spatial covariance matrix $\Sigma$, however, a minimal necessity requirement can also be stated in terms of the following combined matching criterion (Lindeberg 2025b Equation (118)):
@@
3.5.2 Joint spatio-temporal covariance property
@@
  • $u=(u_{1},u_{2})\in{\mathbb{R}}^{2}$ is a 2-D velocity vector that may correspond to the orthographic projection of the 3-D motion field $U$ of the local surface patch observed by the vision system.
@@
The geometric motivation for using this model is beyond the reasons for using a combination of uniform spatial scaling transformations and spatial affine transformations according to the previously stated motivation in Section 3.5.1, that (see Figure 4 for an illustration):
@@
  • the temporal scaling factor $S_{t}$ models the effect of observing similar spatio-temporal events that may occur either faster or slower relative to a previously observed reference view.
@@
provided that the purely spatial parameters $s$, $s^{\prime}$, $A$ and $A^{\prime}$ are matched according to (29) and (30) and additionally the temporal scale parameters $\tau$ and $\tau^{\prime}$ as well as the velocity parameters $v$ and $v^{\prime}$ are matched according to Lindeberg (2025b) Equations (254)–(255):
@@ 3.6 Motivation for filter banks of receptive field responses
A main consequence of the above covariance properties is therefore that from a requirement of the vision system to be covariant under the basic classes of geometric image transformations, if we want to have the ability to match the effects of the either purely spatial or joint spatio-temporal smoothing operations that occur in the models for the receptive fields, then we have to have the ability to make use of receptive field responses computed for matching sets of filter parameters, either $(s,\Sigma)$ or $(s^{\prime},\Sigma^{\prime})$ in the purely spatial case or $(s,\Sigma,\tau,v)$ and $(s^{\prime},\Sigma^{\prime},\tau^{\prime},v^{\prime})$ in the joint spatio-temporal case.
@@ 4 Infinitesimal hybrid Lie semi-group structures for the generalized Gaussian derivative model
Given the above formulations of models for either purely spatial or joint spatio-temporal receptive fields in terms of either purely spatial or joint spatio-temporal derivative expressions applied to either purely spatial or joint spatio-temporal smoothing kernels as obtained from multi-parameter scale-space representations, one may ask how such receptive field responses are related between different values of the filter parameters. Specifically, one may ask if and then how a receptive field response at a coarser level of spatial and temporal scales may be computed from receptive field responses at finer spatial and temporal scales, as illustrated in Figure 5.
Figure 5: The main subject in this paper concerns how to relate receptive field responses between different values of the parameters of the receptive fields, either in terms of (i) infinitesimal relationships closely related to the notions of Lie groups and infinitesimal generators for semi-groups, and (ii) macroscopic relationships in terms of cascade smoothing properties, with close relations to the notion of Lie algebras as can be related to Lie groups via the exponential map. Compared to regular Lie groups and Lie algebras, some of the evolution relations do, however, have a directional preference, implying that the evolution can only be performed in a single direction, and not in the reverse direction.
@@
To a major extent, this will be structurally similar to the notion of a Lie group structure over the filter parameters as expressed in terms of relations between derivatives with respect to the filter parameters $(s,\Sigma)$ or $(s,\Sigma,\tau,v)$ in relation to derivatives with respect to the either spatial or joint spatio-temporal image coordinates $x$ or $(x,t)$. Since the evolution properties with respect to the spatial scale parameter $s$ and the temporal scale parameter $\tau$ can, however, only performed in the direction of increasing values of the scale parameters, the corresponding transformations will therefore not form a full group, but only a semi-group with regard to these two dimensions of variabilities in the filter parameters. For the derived infinitesimal evolution equations with respect to these scale parameters, the resulting notions will therefore be closer to the notion of infinitesimal generators for semi-groups over continuous evolution parameters. For two of the dimensions of the variability, concerning the mixed element $\Sigma_{12}$ in the spatial covariance matrix $\Sigma$ as well as the variabilities with respect to the elements $v_{1}$ and $v_{2}$ in the velocity vector $v$, the variabilities are, on the other hand, not as constrained. For these reasons, we will refer to the resulting structures to be derived as hybrid Lie semi-group structures.
@@ 4.1 Methodology to be used
In this section, we will derive corresponding generalizations of Hermite polynomials for the 2-D and 2+1-D Gaussian kernels that correspond to the either purely spatial or joint spatio-temporal smoothing operations in the spatial and spatio-temporal receptive field models according to (7) and (8).
@@
Stated more formally in a compact manner. Let
@@
then it follows from the linearity of the scale-space representation $L$ that
@@
Specifically, if we let the operator ${\cal D}_{p}$ denote the either purely spatial differentiation operator ${\cal D}_{x}$ used for computing a spatial-derivative-based purely spatial receptive field response according to (7) or alternatively using the same notation ${\cal D}_{p}$ for denoting a spatio-temporal-derivative-based joint spatio-temporal receptive field response according to (8), such that those receptive field responses can be combined on the compact form
@@
then it follows from the linearity of the involved differentiation operators that the infinitesimal relationships between differentiation with respect to a filter parameter $\partial_{P_{i}}$ and the corresponding either purely spatial or joint spatio-temporal differentiation operator ${\cal D}_{i,p}$ according to (57) generalizes to corresponding receptive field responses according to

### v2

@@ 3.3 Visualizations of spatial and spatio-temporal receptive fields
Figures 13 show sample illustrations of receptive fields according to the generalized Gaussian derivative model for visual receptive fields. Figure 1 shows purely spatial receptive fields of the form
$T_{\varphi^{m}}(x;\;s,\Sigma)=\partial_{\varphi^{m}}g(x;\;s,\Sigma),$ (15)
@@
$\begin{split}\lambda_{2}&=\sigma_{2}^{2}.\end{split}$ (20)
This then leads to the following explicit expression for the affine Gaussian derivative kernel (Lindeberg 2024 Equations (163) and (164))
@@ 3.4 Relations to scale-normalized derivatives
In Lindeberg (2021; 2025c; 2025d), so-called unnormalized spatio-temporal derivative expressions of the above forms are additionally combined with different forms of scale normalization, as depending on the actual values of the spatial scale parameter $s$, the spatial covariance matrix $\Sigma$ and the temporal scale parameter $\tau$, to define scale-normalized spatial and spatio-temporal derivative operators [^2] For in-depth treatments about how to extend the regular unnormalized spatial and spatio-temporal derivative operators ${\cal D}_{x}$ and ${\cal D}_{x,t}$ to corresponding scale-normalized spatial and spatio-temporal derivative operators, that make the full derivative-based receptive field responses in (7) and (8) provably covariant under the composed geometric image transformations in (26), (33), (34) and (35), see Lindeberg (2025c; 2025d). Basically, to achieve covariance under with respect to uniform spatial scaling transformations of the form $x^{\prime}=S_{x}\,x$, where $S_{x}\in{\mathbb{R}}_{+}$ is a spatial scaling factor, the partial derivative operators $\partial_{x_{i}}$ in the spatial gradient operator $\nabla_{x}$ according to (10) and the spatial Hessian operator ${\cal H}_{x}$ according to (11) should be replaced by regular scale-normalized spatial derivatives according to $\partial_{x_{i},\text{norm}}=s^{1/2}\,\partial_{x_{i}}$ according to Lindeberg (1998) Equation (6) for the scale normalization power $\gamma=1$, while the regular directional derivative operator $\partial_{\varphi}^{m}$ according to (9) should be replaced by the scale-normalized directional derivative operator $\partial_{\varphi,\text{norm}}^{m}$ according to Lindeberg (2025c) Equation (33). To achieve covariance with respect to non-isotropic spatial affine transformations of the form $x^{\prime}=A\,x$, where $A$ is a positive definite $2\times 2$ affine transformation matrix, the regular gradient operator $\nabla_{x}$ according to (10) should be replaced by the scale-normalized affine gradient operator $\nabla_{x,\text{affnorm}}$ according to Lindeberg (2025c) Equation (111), while the regular Hessian operator ${\cal H}_{x}$ according to (11) should be replaced by the scale-normalized affine Hessian operator ${\cal H}_{x,\text{affnorm}}$ according to Lindeberg (2025c) Equation (140). To achieve both covariance under Galilean transformations of the form $x^{\prime}=x+u\,t$, where $u=(u_{1},u_{2})^{T}\in{\mathbb{R}}^{2}$ is a 2-D motion vector, and covariance under temporal scaling transformations of the form $t^{\prime}=S_{t}\,t$, where $S_{t}\in{\mathbb{R}}_{+}$ is a temporal scaling factor, the temporal differentiation of the receptive fields should be performed in terms of scale-normalized velocity-adapted temporal derivative operators obtained by replacing the regular velocity-adapted temporal derivative operator $\partial_{\bar{t}}^{n}$ in (12) by its scale-normalized correspondence $\partial_{\bar{t},\text{norm}}^{n}$ according to Lindeberg (2025c) Equation (168). [^3] If using scale-normalized derivative operators in the models for the visual receptive fields, it should, however, be observed that the introduction of scale-normalized derivatives will substantially influence the evolution properties over scales of the receptive field responses. This follows, since the scale-dependent scale normalization factors used in the formulations of scale-normalized derivative operators, as extensively described in Lindeberg (2025c), would then also need to be differentiated with respect to the spatial and the temporal scale parameters. For this reason, we do in this paper isolate the evolution properties over the filter parameters to only receptive field models expressed in terms of regular, not scale-normalized, spatial and/or spatio-temporal derivative expressions. If the use of scale-normalized derivatives is warranted, as it can be because of the requirement of making it possible to match the output of derivative-based receptive field operators on the image data, this can be easily accomplished by performing the scale normalization in a second stage, after the computation of regular, not scale-normalized, derivative operators in a first processing stage. that are provably scale covariant, in the sense that the magnitude of the filter responses can be perfectly matched under specific families of geometric image transformations. In this treatment, we will, however, not consider the effects of such scale normalization, and will instead focus solely on the effects of the either purely spatial or joint spatio-temporal smoothing operations, obtained by convolution with kernels of the forms (1) or (2) in combination with either purely spatial or joint spatio-temporal derivative computations, in turn, leading to receptive field responses of the forms (7) or (8). For the task of relating the responses of linear receptive fields of the forms (7) or (8), such an analysis is sufficient, provided that the spatial differentiation operators ${\cal D}_{x}$ and the spatio-temporal differential operators ${\cal D}_{x,t}$ in these receptive field models are purely linear operators.
@@ 3.5 Covariance properties of the spatial and the spatio-temporal smoothing kernels
With regard to variabilities of the shapes of the spatial or the spatio-temporal receptive fields in relation to the variabilities in spatial or spatio-temporal image transformations involved in the image formation process, that determines the structures of the either purely spatial or joint spatio-temporal image data $f$, an important consequence of the formulation of the spatial smoothing kernel $T(x;\;s,\Sigma)$ according to (1) and the formulation of the joint spatio-temporal smoothing kernel $T(x,t;\;s,\Sigma,\tau,v)$ according to (2) is that these forms of convolution kernels are covariant under important basic classes of geometric image transformations.
Figure 4: Illustration of the geometry underlying the composed locally linearized projection models in Equation (26) and Equations (33)–(35), when extended to a multi-view imaging situation, with each view indexed by $k$. We consider a local, in the spatio-temporal case possibly moving, surface patch, which is projected to an arbitrary view in a multi-view locally linearized projection model, with the fixation point $F$ on the surface mapped to the origin $O^{(k)}=0$ in the image plane for the observer with the optic center $P^{(k)}$. Then, any point in the tangent plane to the surface at the fixation point, as parameterized by the local coordinates $\xi$ in a coordinate frame attached to the tangent plane of the surface with $\xi=0$ at the fixation point $F$, is by the local linearization mapped to the image point $x^{(k)}$. (Figure reproduced from Lindeberg (2025c) with permission (OpenAccess).)
3.5.1 Purely spatial covariance properties
@@
$f^{\prime}(x^{\prime})=f(x)\quad\quad\mbox{for}\quad\quad x^{\prime}=S_{x}\,A,$ (26)
@@
provided that the parameters of the receptive fields are matched according to Lindeberg (2025c) Equations (252)–(253):
@@
Due to the overparameterization of the degrees of freedom spanned by the spatial scale parameter $s$ and the spatial covariance matrix $\Sigma$, however, a minimal necessity requirement can also be stated in terms of the following combined matching criterion (Lindeberg 2025c Equation (118)):
@@
3.5.2 Joint spatio-temporal covariance properties
@@
  • $u=(u_{1},u_{2})^{T}\in{\mathbb{R}}^{2}$ is a 2-D velocity vector, that may correspond to the orthographic projection of the 3-D motion field $U$ of the local surface patch observed by the vision system, and
@@
The geometric motivation for using this model is, beyond the reasons for using a combination of uniform spatial scaling transformations and spatial affine transformations according to the previously stated motivation in Section 3.5.1, that (see Figure 4 for an illustration):
@@
  • the temporal scaling factor $S_{t}$ models the effect of observing similar spatio-temporal events, that may occur either faster or slower relative to a previously observed reference view.
@@
provided that the purely spatial parameters $s$, $s^{\prime}$, $A$ and $A^{\prime}$ are matched according to (30) and (31) and additionally the temporal scale parameters $\tau$ and $\tau^{\prime}$ as well as the velocity parameters $v$ and $v^{\prime}$ are matched according to Lindeberg (2025c) Equations (254)–(255):
@@ 3.6 Motivation for filter banks of receptive field responses
A main consequence of the above covariance properties is therefore that from a requirement of the vision system to be covariant under the basic classes of geometric image transformations, if we want to have the ability to match the effects of the either purely spatial or joint spatio-temporal smoothing operations, that occur in the models for the receptive fields, then we have to have the ability to make use of receptive field responses computed for matching sets of filter parameters, either $(s,\Sigma)$ or $(s^{\prime},\Sigma^{\prime})$ in the purely spatial case or $(s,\Sigma,\tau,v)$ and $(s^{\prime},\Sigma^{\prime},\tau^{\prime},v^{\prime})$ in the joint spatio-temporal case.
@@ 4 Infinitesimal hybrid Lie semi-group structures for the generalized Gaussian derivative model
Given the above formulations of models for either purely spatial or joint spatio-temporal receptive fields in terms of either purely spatial or joint spatio-temporal derivative expressions applied to either purely spatial or joint spatio-temporal smoothing kernels, as obtained from multi-parameter scale-space representations, one may ask how such receptive field responses are related between different values of the filter parameters. Specifically, one may ask if and then how a receptive field response at a coarser level of spatial and temporal scales may be computed from receptive field responses at finer spatial and temporal scales, as schematically illustrated in Figure 5.
Figure 5: Schematic illustration of how receptive fields for different parameter values and at different image positions can be interrelated to each other, as constituting the main subject of study in this paper. Here, we derive such relationships between receptive field responses obtained for different values of the shape parameters of the receptive fields, either in terms of (i) infinitesimal relationships closely related to the notions of Lie groups and infinitesimal generators for semi-groups, and (ii) macroscopic relationships in terms of cascade smoothing properties, with close relations to the notion of Lie algebras, as can be related to Lie groups via the exponential map. Compared to regular Lie groups and Lie algebras, some of the evolution relations do, however, have a directional preference, implying that the evolution can only be performed in a single direction, and not in the reverse direction. (In this figure, each line represents a connection between two receptive fields for different values of their parameters and/or their spatial positions.)
@@
To a major extent, this will be structurally similar to the notion of a Lie group structure over the filter parameters, as expressed in terms of relations between derivatives with respect to the filter parameters $(s,\Sigma)$ or $(s,\Sigma,\tau,v)$ in relation to derivatives with respect to the either spatial or joint spatio-temporal image coordinates $x$ or $(x,t)$. Since the evolution properties with respect to the spatial scale parameter $s$ and the temporal scale parameter $\tau$ can, however, only performed in the direction of increasing values of the scale parameters, the corresponding transformations will therefore not form a full group, but only a semi-group with regard to these two dimensions of variabilities in the filter parameters. For the derived infinitesimal evolution equations with respect to these scale parameters, the resulting notions will therefore be closer to the notion of infinitesimal generators for semi-groups over continuous evolution parameters. For two of the dimensions of the variability, concerning the mixed element $\Sigma_{12}$ in the spatial covariance matrix $\Sigma$ as well as the variabilities with respect to the elements $v_{1}$ and $v_{2}$ in the velocity vector $v$, the variabilities are, on the other hand, not as constrained. For these reasons, we will refer to the resulting structures, to be derived, as hybrid Lie semi-group structures.
@@ 4.1 Methodology to be used
In this section, we will derive corresponding generalizations of Hermite polynomials for the 2-D and 2+1-D Gaussian kernels, that correspond to the either purely spatial or joint spatio-temporal smoothing operations in the spatial and spatio-temporal receptive field models according to (7) and (8).
@@
Stated more formally in a compact manner, let
@@
then it follows from the linearity of the scale-space representation $L$, that
@@
Specifically, if we let the operator ${\cal D}_{p}$ denote the either purely spatial differentiation operator ${\cal D}_{x}$ used for computing a spatial-derivative-based purely spatial receptive field response according to (7), or alternatively using the same notation ${\cal D}_{p}$ for denoting a spatio-temporal-derivative-based joint spatio-temporal receptive field response ${\cal D}_{x,t}$ according to (8), such that those receptive field responses can be combined on the compact form
@@
then it follows from the linearity of the involved differentiation operators, that the infinitesimal relationships between differentiation with respect to a filter parameter $\partial_{P_{i}}$ and the corresponding either purely spatial or joint spatio-temporal differentiation operator ${\cal D}_{i,p}$ according to (58) generalize to corresponding receptive field responses according to
@@
By dividing the analysis into these three types of receptive field models, we specifically begin with the simpler purely spatial model, in order to then increase the complexity via the isotropic spatio-temporal model towards the fully affine spatio-temporal model, which obeys covariance properties over the largest set of geometric image transformations considered in this study.